A fair die is rolled. The probability that the first 1 occurs at even number of trails is?

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
10 years ago
Hello student,
he probability that the first 1 occurs at even number of trails is
=P(FS)+P(FFFS)+.......
=(5/6)(1/6)+(5/6)3(1/6)+.....
=(5/36)/(1-(25/36))
=5/11
Thanks and Regards
Shaik Aasif
askIITians faculty